OpenStudy (nbarrera):

Help+MEDAL! Evaluate 6+√14+x -√9x when x=2 . Write the answer as a decimal. Round to the nearest hundredth if necessary. A. 5.76 B. 7 C. 14.24 D. -4

OpenStudy (nbarrera):

\[6+\sqrt{14+2}-\sqrt{9*2}\] I set up the equation for more easier.

OpenStudy (love_ranaa):

the square of 14 + 2 is 4 the square of 9 x 2 will be 4.24 4 + 4.2426 is 8.24 + 6 = 14.24

OpenStudy (love_ranaa):

Sorry, I added on the extra 26 on 4.24

OpenStudy (nbarrera):

Thx so much :)

OpenStudy (love_ranaa):

No problem